A random variable X has the following probability distribution:
X P(X)
-2 0.1
-1 k
0 0.2
1 2k
2 0.3
3 3k
Find K and P(X<2).


Answer by htmentor(1343)

The sum of all the probabilities must equal 1.
Sum P(X) = 0.6 + 6k = 1
6k = 0.4
k = 0.4/6 = 0.0667
P(X<2) = 1 - P(X>=2) = 1 - 0.3 - 3*0.0667 = 0.5